You are the incoming inspector for potato chips – you are to ensure that each bag has 16 ounces or more in it. You want your testing to be at the level of significance of 0.05. You pull a sample of 49 bags of chips from a recent truckload. Your sample statistics are:
x-bar (the sample mean) = 15.9 ounces
s ( the sample standard deviation) = 0.35 ounces
(a) what is the null and alternative hypothesis
(b) one or two tailed test ??
(c) what is the critical z value for your test at the 0.05 level of significance??
(d) what is the calculated z value ??
(e) what is your decision about the load of potato chips ??
-- reject ?? -- not-reject ??

Answer by stanbon(75887)

You are the incoming inspector for potato chips – you are to ensure that each bag has 16 ounces or more in it. You want your testing to be at the level of significance of 0.05. You pull a sample of 49 bags of chips from a recent truckload. Your sample statistics are:
x-bar (the sample mean) = 15.9 ounces
s ( the sample standard deviation) = 0.35 ounces
(a) what is the null and alternative hypothesis
Ho: mu = 16 oz (Claim)
Ha: mu < 16 oz
------------------------
(b) one or two tailed test ??
Ans: one
---------------
(c) what is the critical z value for your test at the 0.05 level of significance??
z= -1.645
------------------------
(d) what is the calculated z value ??
z(15.9) = (15.9-16)/[0.35/sqrt(49)]= -0.1*7/0.35= -2
-----------------------------------------
(e) what is your decision about the load of potato chips ??
-- reject ?? -- not-reject ??
since -2 is less that -1.645, reject Ho
The mean weight of the bags is statistically less than 16 oz.
============================
Cheers,
Stan H.